# The Latency Price of Threshold Cryptosystem in Blockchains

### v1

@@ 3.3 Cryptography Definitions
Let $t_{\sf sec},t_{\sf rec},n$ with $t_{\sf sec}\leq t_{\sf rec}\leq n$ be natural numbers. We refer to $t_{\sf sec}$ and $t_{\sf rec}$ as the secrecy and reconstruction threshold, respectively. Let $\mathcal{X}$ be a input space. Looking ahead, the input space $\mathcal{X}$ denotes the output space of the underlying Multi-shot Byzantine Broadcast protocol.
A $(n,t_{\sf sec},t_{\sf rec})$-threshold cryptosystem $\sf TC$ is a tuple of PPT algorithms $\sf TC=({\sf Setup},{\sf ShareGen},{\sf Eval},{\sf PEval},\allowbreak{\sf PVer},{\sf Comb})$ defined as follows:
@@
    ${\sf ShareGen}(s)\rightarrow\{{\sf pk}_{i},\llbracket{s}\rrbracket_{i}\}_{i\in[n]}.$ The share generation algorithm takes as input a secret $s\in\mathcal{K}$ and outputs a vector of threshold public keys $\{{\sf pk}_{1},\ldots,{\sf pk}_{n}\}$, and a vector of secret shares $(\llbracket{s}\rrbracket_{1},\ldots,\llbracket{s}\rrbracket_{n})$. The $j$-th party receives ($(\{{\sf pk}_{i}\}_{i\in[n]},\llbracket{s}\rrbracket_{j})$.
@@
    ${\sf Comb}(S,{\sf val},\{({\sf pk}_{i},\sigma_{i})\}_{i\in S})\rightarrow\sigma/\bot.$ The combine algorithm takes as input a set $S\subseteq[n]$ with $|S|\geq t_{\sf rec}$, an value ${\sf val}$, and a set of tuples $({\sf pk}_{i},\sigma_{i})$ of public keys and $\sf TC$ output shares of parties in $S$. It outputs either a $\sf TC$ output $\sigma$ or $\bot$.
@@
We require a threshold cryptosystem to satisfy the following Robustness and Secrecy properties.
[…]
Game 3.1 (Robustness Under Chosen Message Attack)
[…]
    1. 1.
    2. 2.
Definition 8 (Robustness Under Chosen Message Attack)
[…]
Next, we describe the secrecy property. Intuitively, the secrecy property ensures that ${\sf Eval}(s,{\sf val})$ remains hidden from an adversary $\mathcal{A}$ that corrupts up to $t_{\sf sec}$ parties, where the precise notion of “hidden” depends on the application. For example, when $\sf TC$ is a threshold signature scheme, i.e., ${\sf Eval}(s,{\sf val})$ is a signature on the message ${\sf val}$ using signing key $s$, the secrecy property is analogous to the standard unforgeability property of signature schemes. Similarly, when $\sf TC$ is a threshold decryption scheme, i.e., ${\sf Eval}(s,{\sf val})$ outputs the decryption of the ciphertext ${\sf val}$ using secret key $s$, the Secrecy property requires that $\sf TC$ is semantically secure in the presence of an attacker $\mathcal{A}$ that corrupts up to $t$ parties.
[…]
Game 3.2 (Unpredictability Under Chosen Message Attack)
[…]
Definition 9 (Unpredictability Under Chosen Message Attack)
[…]
We define global reconstruction and local reconstruction time as follows.
@@
For any $(n,t_{\sf sec},t_{\sf rec})$-threshold cryptosystem with secret $s$ and input ${\sf val}$, we say that ${\sf Eval}(s,{\sf val})$ is globally reconstructed if and only if the adversary learns ${\sf Eval}(s,{\sf val})$ (or equivalently $t_{\sf rec}-|\mathcal{C}|$ honest parties reveal $\sf TC$ output shares to $\mathcal{A}$). The global reconstruction time ${\sf GRT}$ is defined to be the earliest physical time when ${\sf Eval}(s,{\sf val})$ is globally reconstructed.
@@
Double sharing of the secret. Looking ahead, we require our threshold cryptosystem to support double sharing of the same secret for two sets of thresholds $(t_{\sf sec},t_{\sf rec})$ and $(t_{\sf sec}^{\prime},t_{\sf rec}^{\prime})$ where $(t_{\sf sec},t_{\sf rec})\neq(t_{\sf sec}^{\prime},t_{\sf rec}^{\prime})$. It is easy to see that threshold cryptosystems based on Shamir secret sharing [41] easily support double sharing.
@@ 4 Blockchain-Native Threshold Cryptosystem
We now formally define the problem of blockchain-native threshold cryptosystem. As mentioned in Section 1, in such a system, a secret is shared among the parties participating in the blockchain protocol and these parties seek to collaboratively run a threshold cryptographic protocol, after every block, using the shared secret and the block as input.
@@
Let $\sf MBB$ be a Multi-shot Byzantine Broadcast protocol as in Definition 1. Let $\sf TC=({\sf Setup},{\sf ShareGen},{\sf Eval},{\sf PEval},\allowbreak{\sf PVer},{\sf Comb})$ be a $(n,t_{\sf sec},t_{\sf rec})$-threshold cryptosystem as in Definition 6. A blockchain-native threshold cryptosystem protocol $\sf BTC=(\sf MBB,\sf TC)$ is defined as follows.
@@
    Upon $\sf MBB$ outputs ${\sf finalize}(r,m)$ for any round $r\in\mathbb{N}$, parties run the $\sf TC$ protocol to compute $\sigma={\sf Eval}(s,(r,m))$  [^2] More generally, $m$ can be the blockchain state after round $r$. and outputs $(r,m,\sigma)$ [^3] In the case $\sf MBB$ outputs $m=\bot$, computing $\sigma={\sf Eval}(s,(r,m))$ may not be meaningful, and the solutions in the paper can be easily adapted to output $\sigma=\bot$ instead. For the ease of the presentation, we still let parties compute $\sigma={\sf Eval}(s,(r,m))$ for $m=\bot$.
We require $\sf BTC$ to satisfy the following properties, except for negligible probabilities.
@@
  • Termination. After GST, for any round $r\in\mathbb{N}$ each honest party eventually outputs $(r,m,\sigma)$ where $m\in{\mathcal{M}}\cup\{\bot\}$.
@@
  • Secrecy. If an honest party outputs $(r,m,\sigma)$, then $\sigma={\sf Eval}(s,(r,m))$, and the adversary cannot compute ${\sf Eval}(s,(r,m^{\prime}))$ for $m^{\prime}\neq m$.
Example of blockchain-native threshold cryptosystem. To facilitate understanding, as a specific example of blockchain-native threshold cryptosystem, consider blockchain-native distributed randomness that generates a shared randomness for every finalized block. $\sf TC$ for this application can be a threshold VRF scheme. Upon $\sf MBB$ (the blockchain consensus layer) outputs $m$ (a block) for a round $r$, parties run the $\sf TC$ protocol to compute the shared randomness ${\sf Eval}(s,(r,m))$.
For $\sf BTC=(\sf MBB,\sf TC)$, we use ${\sf GRT}_{r}$ to denote the global reconstruction time of round $r$ in $\sf TC$, where the global reconstruction time in $\sf TC$ is defined in Definition 10.
Below we define the latency of a blockchain-native threshold cryptosystem to be the time from its finalization to reconstruction, which measures the latency overhead added by blockchain-native threshold cryptosystem. We define the latency for the feasibility results of this paper. Intuitively, ${\sf L}_{r}$ is the maximum time difference, across all honest parties, between the time a honest party $i$ finalizes $(m,r)$ from $\sf MBB$  and the time the same honest party $i$ outputs $(m,r,{\sf Eval}(s,(m,r)))$.
@@
During an execution $\mathcal{E}$ of a blockchain-native threshold cryptosystem $\sf BTC=(\sf MBB,\sf TC)$, for any round $r$ and party $i$, let $T_{i,r}^{\sf F}$ be the physical time when party $i$ outputs ${\sf finalize}(r,m)$ for some $m$ in $\sf MBB$, and $T_{i,r}^{\sf O}$ be the physical time when party $i$ outputs $(r,m,\sigma)$ for some $m,\sigma$ in $\sf BTC$. The latency of the blockchain-native threshold cryptosystem for round $r$ of execution $\mathcal{E}$ is defined to be ${\sf L}_{r}=\max_{i\in{\mathcal{H}}}(T_{i,r}^{\sf O}-T_{i,r}^{\sf F})$.
In a blockchain-native threshold cryptosystem, such as distributed randomness, the transaction execution relies on the $\sf TC$ output of the threshold cryptosystem. By definition, a party may have to wait a period of ${\sf L}_{r}$ before executing the transactions finalized for round $r$. Therefore, the latency of blockchain-native threshold cryptosystem is directly added to the blockchain’s transaction end-to-end latency.
The following lemma is a direct implication of the blockchain-native threshold cryptosystem definition. Intuitively, a blockchain-native threshold cryptosystem should not reconstruct the $\sf TC$ output before a message is finalized.
Lemma 1
Proof
@@ 4.1 Blockchain-Native Threshold Cryptosystem with Finalization Threshold
A blockchain-native threshold cryptosystem with finalization threshold, or $\sf BTC_{FT}=(\sf MBB_{FT},\sf TC)$, is a blockchain-native threshold cryptosystem (Definition 11) that uses a $\sf MBB_{FT}$ protocol (Definition 5).
For brevity, we will henceforth use $\sf BTC$ to refer to blockchain-native threshold cryptosystem, and $\sf BTC_{FT}$ to refer to blockchain-native threshold cryptosystem with finalization threshold.
As some of the results in the paper hold under optimistic conditions, we formally define these optimistic conditions next.
Definition 14 (Error-free Execution)
We define an execution of a $\sf BTC$ protocol as error-free if all parties are honest.
Definition 15 (Synchronous Execution)
We define an execution of a protocol $\sf BTC_{FT}=(\sf MBB_{FT},\sf TC)$ as synchronous for a round $r$ if all honest parties prefinalize the same message $m\in{\mathcal{M}}$ for $r$ at the same physical time in $\sf MBB_{FT}$.
Definition 16 (Optimistic Execution)
We define an execution of a $\sf BTC_{FT}$ protocol as optimistic if the execution is error-free and synchronous for any round $r$, and all messages have the same delay.
@@ 4.2 A Strawman Protocol
3:let ($(\{{\sf pk}_{j}\}_{j\in[n]},\llbracket{s}\rrbracket_{i})\leftarrow{\sf ShareGen}(s)$ for thresholds $t+1\leq t_{\sf sec}\leq t_{\sf rec}\leq n-t$
@@
2:    let $\bm{m}[r]\leftarrow m$
@@
4:    let $\sigma_{i}\leftarrow{\sf PEval}(\llbracket{s}\rrbracket_{i},(r,m))$
5:    send $({\tt SHARE},r,m,\sigma_{i})$ to all parties
@@
As a warm-up, we first describe a strawman protocol for any blockchain-native threshold cryptosystem $\sf BTC=(\sf MBB,\sf TC)$. The protocol has a latency ${\sf L}_{r}\geq\delta$ for any round $r\in\mathbb{N}$ even in executions that are error-free and have the same message delay $\delta$ between honest parties. To the best of our knowledge, all existing blockchain-native threshold cryptosystem follow this approach. For example, Dfinity [31] blockchain adopts this approach for generating distributed randomness, and thus incurs a latency of at least one message delay. We present the strawman protocol in Algorithm 1, which works for both tight and ramp thresholds. For brevity, we refer to this protocol as the slow path.
@@
As part of the setup phase, each party $i$ receives ($(\{{\sf pk}\}_{i\in[n]},\llbracket{s}\rrbracket_{i})$, where $\llbracket{s}\rrbracket_{i}$ is the secret share of party $i$ and $\{{\sf pk}\}_{i\in[n]}$ is the vector of threshold public keys of all parties. Each party maintains a FIFO $queue$ to record the finalized rounds awaiting the $\sf TC$ output. These rounds are pushed into the FIFO $queue$ in the order they are finalized, and only the top of the FIFO $queue$ will pop and be output. Looking ahead, this ensures Total Ordering, even when parties reconstruct $\sf TC$ outputs of different rounds in out of order. Each party additionally maintains two maps $\bm{m}$ and $\bm{\sigma}$ to store the finalized message and $\sf TC$ output shares of each parties of each round, respectively.
@@
In the strawman protocol, for any given round $r$, each party $i$ waits until a message $m$ is finalized by the $\sf MBB$ protocol in round $r$. Upon finalization, each party $i$ computes its $\sf TC$ output share $\sigma_{i}:={\sf PEval}(\llbracket{s}\rrbracket_{i},(r,m))$ and sends the SHARE message $({\tt SHARE},r,m,\sigma_{i})$ to all parties. Party $i$ also adds round $r$ to $queue$ and updates $\bm{m}$ as $\bm{m}[r]:=m$. Next, upon receiving $({\tt SHARE},r,m,\sigma_{j})$ from party $j$, party $i$ first validates $\sigma_{j}$ using ${\sf PVer}$ algorithm and adds $\sigma_{j}$ to the set $S_{r,m}$ upon successful validation. Finally, upon receiving $t_{\sf rec}$ valid SHARE messages for $(m,r)$, party $i$ computes the $\sf TC$ output $\sigma$ using ${\sf Comb}$ algorithm and updates $\bm{\sigma}$ as $\bm{\sigma}[r]=\sigma$. Whenever party $i$ has the $\sf TC$ output of round $r$ that is the top of $queue$, party $i$ pops the queue and outputs the result $(r,\bm{m}[r],\bm{\sigma}[r])$ for round $r$.
@@
To ensure the Termination property in the strawman protocol, the reconstruction threshold must be greater than $n-t$, i.e., $t_{\sf rec}\leq n-t$. Intuitively, this ensures that once the $\sf MBB$ outputs in a round, every honest party receives a sufficient number of $\sf TC$ output shares to reconstruct the $\sf TC$ output. Additionally, for Secrecy for the strawman protocol, the secrecy threshold must be greater than the number of $\sf TC$ shares controlled by the adversary, i.e., $t_{\sf sec}\geq t+1$. Intuitively, this prevents the adversary from reconstructing the $\sf TC$ output on its own.
Analysis of the strawman protocol. The correctness of the protocol is straightforward and is omitted here for brevity.
@@
We will now argue that the slow path has a latency overhead of at least $\delta$ even in executions that are error-free and have the same message delay $\delta$ between honest parties. More specifically, for any round $r$, any party needs to receive at least $t_{\sf rec}-|\mathcal{C}|\geq 1$ shares from the honest parties to compute $\sigma$. Consider the first honest party $i$ that outputs ${\sf finalize}(r,m)$. In the strawman protocol, party $i$ needs to wait for at least one message delay starting from finalization to receive the shares from the honest parties to compute $\sigma$, since other honest parties only send shares after finalization. This means ${\sf L}_{r}\geq\delta$. As mentioned in Section 1, the latency of $\sf MBB$ can be as short as three message delays, so adding an additional message delay to the system represents a significant overhead.
5 Blockchain-Native Threshold Cryptosystem with Tight Thresholds
Algorithm 2 Blockchain-Native Threshold Cryptosystem with Tight Thresholds
SETUP:
3:let ($(\{{\sf pk}_{j}\}_{j\in[n]},\llbracket{s}\rrbracket_{i})\leftarrow{\sf ShareGen}(s)$ for thresholds $t_{\sf sec}=t_{\sf rec}=t_{\sf fin}$triangleright $t_{\sf fin}$ is the finalization threshold of $\sf MBB_{FT}$

### v3

@@ 2.2 Cryptography Definitions
Let $t_{\sf sec},t_{\sf rec},n$ with $t_{\sf sec}\leq t_{\sf rec}\leq n$ be natural numbers. We refer to $t_{\sf sec}$ and $t_{\sf rec}$ as the secrecy and reconstruction threshold. Let $\mathcal{X}$ be a input space. Looking ahead, the input space $\mathcal{X}$ denotes the output space of the underlying Multi-shot Byzantine Broadcast protocol.
A $(n,t_{\sf sec},t_{\sf rec})$-threshold cryptosystem $\sf TC$ is a tuple of PPT algorithms $\sf TC=({\sf Setup},{\sf ShareGen},{\sf Eval},{\sf PEval},\allowbreak{\sf PVer},{\sf Comb},{\sf Verify})$ defined as follows:
@@
    ${\sf ShareGen}(s)\rightarrow\{{\sf pk},{\sf pk}_{i},\llbracket{s}\rrbracket_{i}\}_{i\in[n]}.$ The share generation algorithm takes as input a secret $s\in\mathcal{K}$ from a secret key space $\mathcal{K}$ and outputs a public key ${\sf pk}$, a vector of threshold public keys $\{{\sf pk}_{1},\ldots,{\sf pk}_{n}\}$, and a vector of secret shares $(\llbracket{s}\rrbracket_{1},\ldots,\llbracket{s}\rrbracket_{n})$. The $j$-th party receives ($(\{{\sf pk}_{i}\}_{i\in[n]},\llbracket{s}\rrbracket_{j})$.
@@
    ${\sf Comb}(S,{\sf val},\{({\sf pk}_{i},\sigma_{i})\}_{i\in S})\rightarrow\sigma/\bot.$ The combine algorithm takes as input a set $S\subseteq[n]$ with $|S|\geq t_{\sf rec}$, a value ${\sf val}$, and a set of tuples $({\sf pk}_{i},\sigma_{i})$ of public keys and $\sf TC$ output shares of parties in $S$. It outputs a $\sf TC$ output $\sigma$ or $\bot$.
  7. 7.
    ${\sf Verify}({\sf pk},{\sf val},\sigma)\rightarrow 0/1:$ The verification algorithm takes as input a public key ${\sf pk}$, input ${\sf val}$, and evaluation output $\sigma$. It outputs 1 (accept) or 0 (reject).
We require a threshold cryptosystem to satisfy the standard Robustness and Secrecy properties, as defined in Section 0.A.1 due to space constraints.
@@
For any $(n,t_{\sf sec},t_{\sf rec})$-threshold cryptosystem with secret $s$ and input ${\sf val}$, we say ${\sf Eval}(s,{\sf val})$ is globally reconstructed if and only if the adversary learns ${\sf Eval}(s,{\sf val})$ (or equivalently $t_{\sf rec}-|\mathcal{C}|$ honest parties reveal $\sf TC$ output shares to $\mathcal{A}$). The global reconstruction time ${\sf GRT}$ is defined to be the earliest physical time when ${\sf Eval}(s,{\sf val})$ is globally reconstructed (i.e., when $t_{\sf rec}-|\mathcal{C}|$ honest parties reveal $\sf TC$ output shares to $\mathcal{A}$).
@@
Double sharing of the secret. Looking ahead, we require our threshold cryptosystem to support double sharing of the same secret for two sets of thresholds $(t_{\sf sec},t_{\sf rec})$ and $(t_{\sf sec}^{\prime},t_{\sf rec}^{\prime})$ where $(t_{\sf sec},t_{\sf rec})\neq(t_{\sf sec}^{\prime},t_{\sf rec}^{\prime})$. Threshold cryptosystems based on Shamir secret sharing [41] easily support double sharing.
@@ 3 Blockchain-Native Threshold Cryptosystem
We now formally define the problem of blockchain-native threshold cryptosystem. In such a system, a secret is shared among the participants in the blockchain protocol and these parties seek to collaboratively run a threshold cryptographic protocol, after every block, using the shared secret and the block as input.
@@
Let $\sf MBB$ be a Multi-shot Byzantine Broadcast protocol as in Definition 1. Let $\sf TC=({\sf Setup},{\sf ShareGen},{\sf Eval},{\sf PEval},\allowbreak{\sf PVer},{\sf Comb},{\sf Verify})$ be a $(n,t_{\sf sec},t_{\sf rec})$-threshold cryptosystem as in Definition 5. A blockchain-native threshold cryptosystem protocol $\sf BTC=(\sf MBB,\sf TC)$ is defined as follows.
@@
    Upon $\sf MBB$ outputs ${\sf finalize}(r,m)$ for any round $r\in\mathbb{N}$, after some possible delay, parties finish the $\sf TC$ protocol to compute $\sigma={\sf Eval}(s,(r,m))$ and outputs $(r,m,\sigma)$.
We require $\sf BTC$ to satisfy the following except for negligible probabilities.
@@
  • Termination. After GST, for any round $r\in\mathbb{N}$ each honest party $i$ eventually outputs $(r,m_{i},\sigma)$ where $m_{i}\in{\mathcal{M}}\cup\{\bot\}$.
@@
  • Secrecy. If an honest party outputs $(r,m,\sigma)$ where $m\in{\mathcal{M}}\cup\{\bot\}$, then $\sigma={\sf Eval}(s,(r,m))$, and the adversary cannot compute ${\sf Eval}(s,(r,m^{\prime}))$ where $m\in{\mathcal{M}}$ and $m^{\prime}\neq m$.
For $\sf BTC=(\sf MBB,\sf TC)$, we use ${\sf GRT}_{r}$ to denote the global reconstruction time of round $r$ in $\sf TC$, as defined in Definition 7.
Example of blockchain-native threshold cryptosystem.
On-chain distributed randomness generates a shared randomness for every finalized block. $\sf TC$ for this application can be a threshold VRF scheme. Upon $\sf MBB$ (the blockchain consensus layer) outputs $m$ (a block) for a round $r$, parties run the $\sf TC$ protocol to compute the shared randomness ${\sf Eval}(s,(r,m))$.
Below we define the latency of a blockchain-native threshold cryptosystem to measure the introduced latency overhead. Intuitively, ${\sf L}_{r}$ is the maximum time difference, across all honest parties, between the time a honest party $i$ finalizes in $\sf MBB$  and the time the same honest party $i$ outputs. Since the transaction execution relies on the $\sf TC$ output of the threshold cryptosystem, by definition, a party may have to wait a period of ${\sf L}_{r}$ before executing the transactions finalized for round $r$, thus increasing the blockchain’s transaction end-to-end latency.
@@
During an execution of a blockchain-native threshold cryptosystem $\sf BTC=(\sf MBB,\sf TC)$, for any round $r$ and party $i$, let $T_{i,r}^{\sf F}$ be the physical time when party $i$ outputs ${\sf finalize}(r,m)$ for some $m$ in $\sf MBB$, and $T_{i,r}^{\sf O}$ be the physical time when party $i$ outputs $(r,m,\sigma)$ for some $m,\sigma$ in $\sf BTC$. The latency for round $r$ of the execution is defined to be ${\sf L}_{r}=\max_{i\in{\mathcal{H}}}(T_{i,r}^{\sf O}-T_{i,r}^{\sf F})$.
@@ 3.1 Blockchain-Native Threshold Cryptosystem with Finalization Threshold
A blockchain-native threshold cryptosystem with finalization threshold, or $\sf BTC_{FT}=(\sf MBB_{FT},\sf TC)$, is a blockchain-native threshold cryptosystem (Definition 8) that uses an $\sf MBB_{FT}$ protocol (Definition 4).
We will henceforth shorten blockchain-native threshold cryptosystem as $\sf BTC$, and blockchain-native threshold cryptosystem with finalization threshold as $\sf BTC_{FT}$.
Some of the paper’s results hold under optimistic conditions defined below.
Error-free. An execution of $\sf BTC$ is error-free if all parties are honest.
Synchronized execution. An execution of $\sf BTC_{FT}=(\sf MBB_{FT},\sf TC)$ is synchronized for a round $r$ if all honest parties prefinalize the same message $m\in{\mathcal{M}}$ for $r$ at the same physical time  [^2] This condition is defined solely for proving theoretical latency claims. In practice, different honest parties may prefinalize at different physical times, and the result of the paper still achieves latency improvements as in Section 6.1. in $\sf MBB_{FT}$.
Optimistic. An execution of $\sf BTC_{FT}$ is optimistic if the execution is error-free and synchronized for any round $r$, and all messages have the same delay.
@@ 3.2 A Strawman Protocol
3:let ($(\{{\sf pk}_{j}\}_{j\in[n]},\llbracket{s}\rrbracket_{i})\leftarrow{\sf ShareGen}(s)$ for thresholds $t+1\leq t_{\sf sec}\leq t_{\sf rec}\leq n-t$triangleright $s$ is unknown to any party
@@
2:    let $\bm{m}[r]\leftarrow m$ and $\sigma_{i}\leftarrow{\sf PEval}(\llbracket{s}\rrbracket_{i},(r,m))$
@@
4:    send $({\tt SHARE},r,m,\sigma_{i})$ to all parties
@@
As a warm-up, we first describe a strawman protocol for any blockchain-native threshold cryptosystem $\sf BTC=(\sf MBB,\sf TC)$ (Definition 8) in Algorithm 1, which works for both tight and ramp thresholds. The protocol has a latency ${\sf L}_{r}\geq\delta$ for any round $r\in\mathbb{N}$ even in error-free executions with constant message delay $\delta$ between honest parties. To the best of our knowledge, all existing blockchain-native threshold cryptosystem follow this approach, such as the Dfinity [31] and Sui [48] blockchains for distributed randomness. For brevity, we refer to this protocol as the slow path.
@@
As part of the setup phase, each party $i$ receives ($(\{{\sf pk}\}_{i\in[n]},\llbracket{s}\rrbracket_{i})$, where $\llbracket{s}\rrbracket_{i}$ is the secret share of party $i$ and $\{{\sf pk}\}_{i\in[n]}$ is the vector of threshold public keys of all parties. Each party maintains a First-in-first-out (FIFO) $queue$ to record the finalized rounds awaiting the $\sf TC$ output. These rounds are pushed into the FIFO $queue$ in the order they are finalized, and only the head of the FIFO $queue$ will pop and be output. Looking ahead, this ensures Total Ordering, even when parties reconstruct $\sf TC$ outputs of different rounds in out of order. Each party additionally maintains two maps $\bm{m}$ and $\bm{\sigma}$ to store the finalized message and $\sf TC$ output shares of each parties of each round, respectively.
@@
In the protocol, for any given round $r$, each party $i$ waits until a message $m$ is finalized by the $\sf MBB$ protocol in round $r$. Upon finalization, each party $i$ computes its $\sf TC$ output share $\sigma_{i}\leftarrow{\sf PEval}(\llbracket{s}\rrbracket_{i},(r,m))$ and sends the SHARE message $({\tt SHARE},r,m,\sigma_{i})$ to all parties. Party $i$ also adds round $r$ to $queue$ and updates $\bm{m}$ as $\bm{m}[r]\leftarrow m$. Next, upon receiving $({\tt SHARE},r,m,\sigma_{j})$ from party $j$, party $i$ first validates $\sigma_{j}$ using ${\sf PVer}$ algorithm and adds $\sigma_{j}$ to the set $S_{r,m}$ upon successful validation. Finally, upon receiving $t_{\sf rec}$ valid SHARE messages for $(m,r)$, party $i$ computes the $\sf TC$ output $\sigma$ using ${\sf Comb}$ algorithm and updates $\bm{\sigma}$ as $\bm{\sigma}[r]=\sigma$. Whenever party $i$ has the $\sf TC$ output of round $r$ that is the head of $queue$, party $i$ pops the queue and outputs the result $(r,\bm{m}[r],\bm{\sigma}[r])$ for round $r$.
@@
To ensure the Termination property, the reconstruction threshold must be no greater than $n-t$, i.e., $t_{\sf rec}\leq n-t$. Intuitively, this ensures that once the $\sf MBB$ outputs in a round, every honest party receives a sufficient number of $\sf TC$ output shares to reconstruct the $\sf TC$ output. Additionally, for Secrecy for the strawman protocol, the secrecy threshold must be greater than the number of $\sf TC$ shares controlled by the adversary, i.e., $t_{\sf sec}\geq t+1$. Intuitively, this prevents the adversary from reconstructing the $\sf TC$ output on its own. The correctness of the protocol is straightforward and is omitted here for brevity.
@@
We will now argue that the slow path has a latency overhead of at least $\delta$ even in error-free executions with constant message delay $\delta$ between honest parties. For any round $r$, any party needs to receive at least $t_{\sf rec}-|\mathcal{C}|\geq 1$ shares from the honest parties to compute $\sigma$. Consider the first honest party $i$ that outputs ${\sf finalize}(r,m)$. In the strawman protocol, party $i$ needs to wait for at least one message delay starting from finalization to receive the shares from the honest parties to compute $\sigma$, since other honest parties only send shares after finalization. Adding one additional message delay to the system represents a significant overhead as the $\sf MBB$ latency can be as short as three message delays
4 Tight Blockchain-Native Threshold Cryptosystem
Algorithm 2 Tight Blockchain-Native Threshold Cryptosystem
SETUP is same as Algorithm 1 except that $t_{\sf sec}=t_{\sf rec}=t_{\sf fin}$.